Use the elimination method to solve the following system of equations. what is x-4y+3z=1
-x+y-5z=-10
2x-4y+3z=-3

Thank you!

Answer by TimothyLamb(4379)

x-4y+3z=1
-x+y-5z=-10
2x-4y+3z=-3
---
put the system of linear equations into standard form
---
x-4y+3z=1
-x+y-5z=-10
2x-4y+3z=-3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -4
y = 1
z = 3
---
now by elimination:
---
x - 4y + 3z = 1
-x + y - 5z = -10
---
add the above two equations:
---
-3y - 2z = -9
---
2(-x + y - 5z = -10)
2x - 4y + 3z = -3
---
-2x + 2y - 10z = -20
2x - 4y + 3z = -3
---
add the above two equations:
---
-2y - 7z = -23
---
-3(-2y - 7z = -23)
2(-3y - 2z = -9)
---
6y + 21z = 69
-6y - 4z = -18
---
add the above two equations:
---
17z = 51
z = 3
---
-3y - 2z = -9
-3y - 2*3 = -9
-3y = -3
y = 1
---
x - 4y + 3z = 1
x = 1 - (- 4y + 3z)
x = 1 + 4y - 3z
x = 1 + 4*1 - 3*3
x = -4
---
solution by elimination:
x = -4
y = 1
z = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---